our sir told that the sum of the interior angles of an 'n' sided polygon is '(n-2)pie'.how?

Grade:12

1 Answers

Anil Pannikar AskiitiansExpert-IITB
13 years ago

Dear Vivek,

 

A quadrilateral has  4 sides, draw a diagonal and divide it into 2 traingles

so , Sum of angle  =  2 * 180 = 360

A pentagon has 5 sides, draw 2 diagonal and divide it into 3 traingles

so , Sum of angle  =  3 * 180 = 540

similarly. a hexagon has 6 sides, draw 3 diagonal and divide it into 4 traingles

so , Sum of angle  =  4* 180 = 720

 

so u can generalize, sum = ( n-2 )*180

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Anil Pannikar

IIT Bombay